# Subword models struggle with word learning, but surprisal hides it

### v1

@@ 6 Discussion and conclusion
This study set out to explore how LMs learn what valid words are, not when words are validly used. We have shown that lexical decision complements surprisal-based approaches and reveals difficulties in word learning that surprisal hides: subword LMs actually struggle with it, whereas character models master this task with ease. In subword LMs, lexical and syntactic learning are inseparable (unlike in humans, see Brinchmann et al., 2019; Donnelly et al., 2025), whereas word learning in character models precedes syntactic learning; the processes are related, yet separable. The a priori token splitting in subword models basically preempts word learning in them, whereas character models pass through more plausible developmental stages. Why this is the case remains open to further inquiry, but it is hard to imagine that arbitrary subword units lead to human-like, cognitively plausible word-level representations (like in exemplar models of lexical storage, cf. Bybee, 2010), whereas character models possibly offer a more justified level of granularity for them. In any case, as we have shown, decisions about the representational levels of LMs tremendously influence their learning pathways.
@@
The generalizability of our findings is constrained by a few factors. The present study has only focused on the English language, but it is plausible that other languages with different writing systems or graphematic and phonotactic rule systems exhibit different patterns under different tokenization schemes. Here, phonetic transcriptions might provide a viable alternative, but real narrow transcriptions that accurately capture the whole breadth of human input are scarce and extremely costly and laborious to manually produce. Besides, for the character LMs, we focus only on small models, as very large models with such tokenization, especially ones providing intermediate checkpoints, are nonexistent at this moment. It would still be interesting to see how they compare to subword models in a setting where parameter size and training data are greatly increased.
From a developmental perspective, word learning in humans also includes other processes than statistical pattern recognition from the input: semantic aspects and real-world reference are equally important, as are multimodal input and communicative intent. The ongoing form-vs.-function debate on LMs Mahowald et al. (2024) has begun to consider these aspects, and further studies should aim at incorporating them.
Finally, we also want to mention that there are attempts to add more linguistic theory to tokenizers. Looking into different tokenizers that, for exameple, try to be more morphology-aware Hofmann et al. (2022); Bauwens and Delobelle (2024) or implement other optimization tricks Schmidt et al. (2024) could yield even more fine-grained points of comparison, but for the present study and its limited scope we focused on the most popular tokenization scheme (BPE) and the most linguistically minimalist alternative (characters).
@@
Due to the nature of this work, no concrete ethical aspects or repercussions need to be discussed. However, we would like to stress that, of course, BabyLMs are not real babies, but only abstractions of a very specific part of their learning capacity (frequency-driven, domain-general learning mechanisms such as entrenchment or resonance), and therefore all claims about their implications for language development in the real world should be interpreted in this light.
@@ Appendix A Model hyperparameters and training details
Our models were trained on a Apple M2 Pro processor with the MPS backend. For the small models, training lasted approx. 20min, for the medium-sized models it took approx. 1h and for the largest models approx. 10h (no principal differences between character and subword models). We share our models on the HuggingFace hub under [link anonymized].
Loss curves for models can be found in Figure 3. For the test loss, we evaluated perplexity on a held-out portion of our training corpus that is comparable in composition. We find no principal differences in loss development, although the character models converge faster. Larger models also tend to converge faster and generally reach smaller absolute loss values. As the similar train and test loss curves indicate, all Llama models succeed in optimizing for their next-token prediction objective. However, it remains questionable how much these scores are constrained by the comparatively small capacity of our BabyLMs.
@@ Appendix B Data creation and tokenization analysis
Word frequency influences lexical decision performance greatly McClelland and Rumelhart (1981); Allen et al. (2005). To incorporate this effect into our study, we create two distinct data sets from words included in wuggy: (i) high-frequency stimuli with a frequency score over 7.0 (occurrences per 1M words in BNC, COCA and other English corpora, as reported in CELEX, Baayen et al., 1995) and (ii) low-frequency stimuli with a frequency score over 0.0 but below 0.7 (so at least one order of magnitude lower). We opted to rely on the CELEX frequency scores because we compare models trained on different corpora – our models are trained on the 10M BabyLM corpus, the models by Goriely et al. (2024) are trained on the 100M BabyLM corpus and the Pythia models are trained on The Pile Gao et al. (2020). As such, frequency scores from these corpora would taint analyses of other models and hinder comparability. For the contextualised stimuli, we sample sentences from the OpenSubtitles Lison and Tiedemann (2016) portion of the BabyLM 2024 corpus Choshen et al. (2024). Both Wuggy and BabyLM data are licensed under the MIT license [^4] As per license information found at https://github.com/WuggyCode/wuggy and https://github.com/babylm/evaluation-pipeline-2024., therefore we release our own stimuli artifacts on HuggingFace under [link anonymized] under the same license. The data contain no information that names or uniquely identifies individual people or offensive content, and are commonly used in computational linguistics.
@@ Appendix F Development of word/non-word differences
In Figure 7, we plot the average difference between word and non-word negative log-probability values across training (for high-frequency data). Positive scores indicate preference for real words. For the character models, the differences are less pronounced and get most extreme at the end of pre-training (where accuracy scores do not change anymore), especially for the lexical decision data, which is already consistent at very early training stages. For the BPE models, we see that at the beginning they actually prefer non-words in the lexical decision task. Only after the first 10% of training they begin to discern words and non-words. Figure 8 shows the equivalent difference curves to Figure 7 for our low-frequency data. While general tendencies remain the same, the absolute differences are generally lower and the differences between the curves are less pronounced.
Figure 7: Average differences between surprisal scores
Figure 8: Average differences between surprisal scores (for low-frequency data)

### v2

@@ 6 Conclusion
We have shown that the lexical decision approach to the study of word learning in LMs complements surprisal-based approaches and reveals difficulties that surprisal hides: subword LMs struggle with lexical decision, whereas character models master this task with ease. Additionally, in subword LMs, lexical and syntactic learning are inseparable, whereas word learning in character models precedes syntactic learning; the processes are related, yet separable. It is plausible that the a priori token splitting in subword models preempts a word discovery process in them, whereas character models first have to pass through this developmental stage, possibly in a somewhat more human-like manner. In any case, as we have shown, decisions about the representational levels of LMs tremendously influence their learning pathways on the different levels of linguistic analysis.
@@
The generalizability of our findings is constrained by a few factors. The present study has only focused on the English language, but it is plausible that other languages with different writing systems or graphematic and phonotactic rule systems exhibit different patterns under different tokenization schemes. Here, phonetic transcriptions might provide a viable alternative, but real narrow transcriptions that accurately capture the whole breadth of human input are scarce and extremely costly and laborious to manually produce (although novel datasets like Goriely and Buttery, 2025b provide an alternative through automatic transcription). Besides, for the character LMs, we focus only on small models, as very large models with such tokenization, especially ones providing intermediate checkpoints, are nonexistent at this moment. It would still be interesting to see how they compare to subword models in a setting where parameter size and training data are greatly increased.
As already mentioned in Section 2, from a developmental perspective, word learning in humans also includes other processes than statistical pattern recognition from the input: semantic aspects and real-world reference are equally important, as are multimodal input and communicative intent. The ongoing form-vs.-function debate on LMs Mahowald et al. (2024) has begun to consider these aspects, and further studies should aim at incorporating them; for example, the object naming paradigm used in many developmental psychology studies would lend itself naturally to the study of multimodal models.
Finally, we also want to mention that there are attempts to add more linguistic theory to tokenizers. Looking into different tokenizers that, for example, try to be more morphology-aware Hofmann et al. (2022); Bauwens and Delobelle (2024); Yehezkel and Pinter (2023); Uzan et al. (2024) or implement other optimization tricks Schmidt et al. (2024) could yield even more fine-grained points of comparison, but for the present study and its limited scope we focused on the most popular tokenization scheme (BPE) and the most linguistically minimalist alternative (characters).
@@
Due to the nature of this work, no concrete ethical aspects or repercussions need to be discussed. However, we would like to stress that, of course, BabyLMs not supposed to simulate real babies, but only abstractions of a very specific part of their learning capacity (frequency-driven, domain-general learning mechanisms such as entrenchment or resonance), and therefore all claims about their implications for language development in the real world should be interpreted in this light.
Acknowledgments
We would like to thank the three anonymous ARR reviewers for their insightful comments and the highly valuable discussions throughout the rebuttal period.
This research has been funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – CRC-1646, project number 512393437, project A02.
@@ Appendix A Model hyperparameters and training details
Our models were trained on a Apple M2 Pro processor with the MPS backend. For the small models, training lasted approx. 20min, for the medium-sized models it took approx. 1h and for the largest models approx. 10h (no principal differences between character and subword models). We share our models with their checkpoints on the HuggingFace hub. [^4] https://huggingface.co/collections/bbunzeck/word-learning-in-small-lms-67bdc218688856162b3be08f
Loss curves for all models can be found in Figure 3. For the test loss, we calculated the perplexity over a held-out portion of our training corpus that is comparable in composition to the training data. We find no principal differences in loss development, although the character models converge faster. Larger models also tend to converge faster and generally reach smaller absolute loss values. As the similar train and test loss curves indicate, all Llama models succeed in optimizing for their next-token prediction objective. However, it remains open to further inquiry how much these scores are constrained by the comparatively small capacity of our BabyLMs, and whether larger models would enhance performance further.
@@ Appendix B Data creation and tokenization analysis
Word frequency influences lexical decision performance greatly McClelland and Rumelhart (1981); Allen et al. (2005). To incorporate this effect into our study, we create two distinct data sets from words included in wuggy: (i) high-frequency stimuli with a frequency score over 7.0 (occurrences per 1M words in BNC, COCA and other English corpora, as reported in CELEX, Baayen et al., 1995) and (ii) low-frequency stimuli with a frequency score over 0.0 but below 0.7 (so at least one order of magnitude lower). We opted to rely on the CELEX frequency scores because we compare models trained on different corpora – our models are trained on the 10M BabyLM corpus, the models by Goriely et al. (2024) are trained on the 100M BabyLM corpus and the Pythia models are trained on The Pile Gao et al. (2020). As such, frequency scores from these corpora would taint analyses of other models and hinder comparability. For the contextualised stimuli, we sample sentences from the OpenSubtitles Lison and Tiedemann (2016) portion of the BabyLM 2024 corpus Choshen et al. (2024). Both Wuggy and BabyLM data are licensed under the MIT license [^5] As per license information found at https://github.com/WuggyCode/wuggy and https://github.com/babylm/evaluation-pipeline-2024., therefore we release our own stimuli artifacts on HuggingFace [^6] https://huggingface.co/datasets/bbunzeck/lexical-decision under the same license. The data contain no information that names or uniquely identifies individual people or offensive content, and are commonly used in computational linguistics.
@@ Appendix F Development of word/non-word differences
In Figure 7, we plot the average difference between word and non-word negative log-probability values across training, for both high-frequency and low-frequency data. Positive scores indicate preference for real words. For the character models, the differences are generally less pronounced and get most extreme at the end of pre-training (where accuracy scores do not change anymore), especially for the lexical decision data, which is already consistent at very early training stages. For the BPE models, we see that at the beginning they actually prefer non-words in the lexical decision task. Only after the first 10% of training they begin to discern words and non-words. While overall tendencies remain the same for both frequency conditions, the absolute differences are generally lower and the differences between the curves are less pronounced in the low-frequency setting.
(a) High-frequency data
(b) Low-frequency data
Figure 7: Average differences between surprisal values across pretraining